The length of a rectangle is twice its width.
If the perimeter of the rectangle is 54, find its area?

Answer by checkley77(12844)

L=2W
LW=54
2W(W)=54
2W^2=54
W^2=54/2
W^2=27
W=SQRT27
W=5.196 ANS. FOR THE WIDTH.
L=2*5.196=10.392 ANS. FOR THE LENGTH.
PROOF:
10.392*5.196=54
54~54